Status Report

ISS On-Orbit Status 5 Aug 2002

By SpaceRef Editor
August 5, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below.

The crew was thanked by the TVIS ground team for the temporary repair
of the treadmill chassis yesterday.  The next procedure for
getting the TVIS back into operation by the crew was promised to be
ready very soon.

CDR Valery Korzun and FE-1 Peggy Whitson began preparations for the
upcoming two Orlan-M spacewalks, EVA-7 and EVA-8, from the DC-1
docking module.  EVA-7 will be on 8/16, EVA-8 on 8/23.
 [Today’s work: locating, identifying and gathering of a long
list of required equipment and tools, such as soft and semi-rigid
transport containers, tethers, storage bags, cable clamps, Velcro,
Aramid and insulating tape, wrenches, wire cutters, wire ties, bungee
cords, hammer, screwdriver, etc.  Later in the day, Korzun and
Whitson had three hours set aside for reviewing and studying EVA-7
procedures and preliminary timeline.]

Also as part of EVA preparations, FE-2 Treschev hooked up the BNP
portable O2 repress tank in the Service Module Transfer Compartment
(PkhO) main oxygen repress line and performed a leak check.
 When filled with O2, the BNP will be installed in the DC-1
docking compartment in support of the Orlan-EVAs.

The crew began another session of the NASA renal (kidney) stone
experiment, with Whitson setting up the necessary hardware and Korzun
beginning his diet log.  [This long-range preventive medicine
investigation features daily random ingestion of either potassium
citrate or placebo tablets.  It is a double-blind JSC research
study investigating methods to prevent formation of kidney stones in
zero-G.  Part of the experiment consists in keeping a metabolic
diet log (food and fluid intake), followed by collection of urine
samples several times per day during each week-long session.].

Korzun performed the mandatory Russian pre-EVA MedOps procedure
MO-6 (hand-cycle ergometry), with FE-2 Sergei Treschev assisting.
[Previous cosmonauts have shown noticeable decrease in arm muscle
tone, and TsUP/IBMP (MCC-Moscow/Institute of Biomedical Problems)
physical fitness experts have made the tolerance hand grip test
analysis (hand ergometry) a standard pre-EVA requirement.  MO-6
today required the CDR to don the ECG (electrocardiogram) biomed
harness, attach three skin electrodes and plug the harness into the
PKO medical exam panel on the cycle ergometer.  The exercise
itself started after 10 seconds of complete rest, by manually
rotating the cycle’s pedals, set at 150 W, backwards until
"complete exhaustion".]

Peggy Whitson initiated battery charging for the upcoming EVA-7,
in particular for the PGTs (pistol grip tools), helmet lights and
REBAs (rechargeable EVA battery assembly).  [To prevent a
repeat of recent battery failures, MCC-H uplinked a slot assignment
list for placing the batteries in the BSA (battery stowage assembly),
which enables the ground to receive battery voltages directly via
telemetry, thus helping to chart the health of each over time.]

Peggy also started recharge of two 120-V batteries (using a
Dual-Cannon Lithium Ion battery charger), in preparation for
tomorrow’s television "production" by the crew in support
of an educational event with the Maryland Science Center.

Tomorrow’s live downlink to the Maryland Science Center (MSC) event
is scheduled for 9:35-9:55 am EDT.  Participants on the ground
include Jim O’Leary (Senior Director of IMAX and Planetarium at the
MSC), Roger Crouch (Senior ISS Scientist at NASA HQ), representatives
from Johns Hopkins University & the Space Telescope Science
Institute, middle school students from three Maryland counties, and
local media.  The downlink will be on NASA-TV.

Sergei Treschev performed daily routine maintenance of the SOSh
environment control and life support system as well as the daily
preparation of the IMS (inventory management system)
"delta" file.  Peggy Whitson did the daily payload
status check.

All crewmembers completed their daily physical exercise program.

After the successful restart of the SM TVM (Terminal Computer system)
on 8/2, all three lanes continue to work in the redundant set.

The Elektron oxygen generator was activated early this morning after
some more troubleshooting and is working at present, but apparently
in a somewhat off-nominal mode.

Last week (8/1), working from the task list, CDR Korzun and FE-1
Whitson took the periodic CMO (crew medical officer) proficiency
training, for half an hour.  [To maintain CMO skills,
crewmembers spend one hour per month reviewing HMS (health
maintenance system) and ACLS (advanced cardio life support) equipment
and procedures via HMS and ACLS CBT (computer-based tŽfýing).
 Impacts of not maintaining this proficiency could lead to a
substantial risk to ISS operations, potential evacuation of the
station, and loss of crew life.]

When after the station reboost on 8/1 an RPC (remote power
controller) controlling CETA (crew and equipment translation aid)
lights opened its circuit, there was some surprise on the ground.
 It now has been determined that the removal of operational CETA
light power during the transition from reboost maneuvers back to
nominal attitude control is a normal feature of the C&C MDM
(command & control computer) software.  [It was designed
for post-EVA transitions to nominal attitude control, with the intent
that the CETA lights would not be needed post-EVA and this automatic
feature would save crew and/or ground command time.]

Today’s targets for the CEO (crew earth observation) program were
Congo-Zimbabwe Biomass Burning (on this pass the best views
of Zimbabwe were to the right of track.  The crew was asked to
try for high oblique views of the smoke buildup in the atmosphere
over this region)
, High Central Andean Glaciers
(afternoon cloudiness over the higher elevations often precludes
photos of this target area.  On this pass visibility was
expected to be better than usual to the left of track.  Crew to
try for context views of the small ice fields and glaciers along the
crest of the Andes)
, Eastern Mediterranean Dust (high
pressure should still be holding over this region.  As ISS
crossed the Sinai, the crew was to look left of track for aerosol
accumulations over the waters of the waters of the eastern
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 12:42 pm

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
 U.S. CDRA CO2 scrubber is Off (leak).  BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify

SM Working Compartment:  Pressure (mmHg) — 761,
temperature (deg C) — 26.7, ppO2 (mmHg) — 157.9, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 761,
temperature (deg C) — 20.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 22.3.
Node:  Pressure (mmHg) — 753.43, temperature (deg C) —
25.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 754.82, temperature (deg C)
— 25.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 754.82,
temperature (deg C) — 29.0; shell heater temp (deg C) — 27.9, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 24.2
PMA-2:  Shell heater temp (deg C) — 13.9

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) were replaced, but some more work
needs to be done).

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in
Autotrack mode.
SM batteries: Data are static.  Battery #1 is off line, battery
#4 is cycling; all other batteries (6) are in "Partial
charge" mode.
FGB: Battery #5 is offline; battery #1 is cycling; all other
batteries (4) are in “Partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs online.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power &
data grapple fixture 1) and MBS PDGF 4, with Keep Alive power on both
strings (based on MBS).
MBS:  Operational on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of this morning, 8:36 am EDT [=
Mean altitude — 395.7 km
Apogee — 407.9 km
Perigee — 383.4 km
Period — 92.5 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0018093
Orbits per 24-hr. day — 15.57
Solar Beta Angle — 34.0 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. “98) — 21172
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.